OpenStudy (anonymous):

can some one help me find the the local extrema for 2x^5-5x^2+1

OpenStudy (amistre64):

take the derivative and find the zeros

OpenStudy (amistre64):

D(2x^5 -5x^2 +1) = 10x^4 -10x 10x^4 = 10x x^4 = x ...... mu guess is at 1 and -1

OpenStudy (amistre64):

x(x^3 - 1) (10)(x)(x-1)(x^2 +x +1) or quite possibly at: x = 0, 1

OpenStudy (amistre64):

check 2nd derivative for concavity...

OpenStudy (amistre64):

or plug these x value sinto the original and see whats bigger :)

OpenStudy (amistre64):

f(0) = 1 f(1) = -2

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=2x%5E5-5x%5E2%2B1 gives you a good picture of it

OpenStudy (anonymous):

awesome thank you very much